OpenStudy (anonymous):

A stone whirled on a string experiences a centripetal acceleration of 10 m/s². If the string were shortened to half its length (one-half the radius) and the speed were doubled, calculate the centripetal acceleration. How can I find this without knowing either speed or the radius?

OpenStudy (anonymous):

@douglaswinslowcooper Do you understand this one?

OpenStudy (anonymous):

a = v^2/r = 10 m/s^2 a' = (2v)^2/(0.5r) a'/a = 8 a' = 80 m/s^2

OpenStudy (anonymous):

Okay thank you very much